OpenStudy (anonymous):

what are the factors of 108? and what is the GCF for108?

OpenStudy (anonymous):

GCF between 108 and what other number?

OpenStudy (anonymous):

54

OpenStudy (anonymous):

greatest COMMON factor assumes two different numbers

OpenStudy (anonymous):

\[108=2^2\times 3^3\] \[54=2\times 3^3\] and so gcf of 54 and 108 is 54 since 54 itself is a factor of 108

OpenStudy (anonymous):

thanks:)